A fruit punch that is 40% alcohol is mixed with a fruit punch that is 30% alcohol. How many gallons of each punch are used to make 5 gallons of punch that is 36% alcohol?
Answer by stanbon(75887)

A fruit punch that is 40% alcohol is mixed with a fruit punch that is 30% alcohol. How many gallons of each punch are used to make 5 gallons of punch that is 36% alcohol?
--------
Equation:
alc + alc = alc
0.40x + 0.30(5-x) = 0.36*5
----
40x + 30*5 - 30x = 36*5
10x = 6*5
x = 3 gallons (amt. of 40% needed)
5-x = 2 gallons( amt. of 30% needed)
=============
Cheers,
Stan H.